One of the angles of a triangle is 8 deg more than 5 times another angle. If the third angle is 4 deg larger than the difference of the first two angles, find the angles.
Found 2 solutions by greenestamps, math_helper:
Answer by greenestamps(13216)



Let the first angle measure be x.

Then the second angle measure is 8 more than 5 times the first: 5x+8.

And the third angle measure is 4 more than the difference between the first two angles: ((5x+8)-x)+4 = 4x+12.

The angles are x, 5x+8, and 4x+12. Write and solve the equation that says the sum of the three angles is 180 degrees.

Then use the value of x to find the measures of the three angles.


Answer by math_helper(2461)


Call the angles A, B, and C.
B = 8+5A
C = (B-A)+4 = ((8+5A)-A)+4 = 4A+12
So we have:
A + (8+5A) + (4A+12) = 180
10A + 20 = 180
10A = 160
A = 16 --> B = 88 and C = 76 (angles in degrees)